Solve the system of equations by elimination.
2x + y = 13
4x + 5y = 11

Answer by ikleyn(52786)

.
2x +  y = 13,     (1)
4x + 5y = 11.     (2)


Multiply eq(1) by the factor 2 (both sides). Keep the equation (2) as is.  You will get an equivalent system


4x + 2y = 26      (1')
4x + 5y = 11      (2')


Now subtract  eq(1') from eq(2').  The terms "4x" will cancel each other,    (<<<---=== it is how the Elimination method works)
and you will get a single equation for the unknown "y"

5y - 2y = 11 - 26,   or

3y = -15,

which implies  y = -15/3 = -5.


Now from equation (1),  2*x - 5 = 13;  hence,  2x = 13 + 5 = 18   and  x = 18/2 = 9.


    Check the solution on your own by substituting the found values into the original equation.

    Do it on your own (I just did it mentally).


Answer  x= 9;  y= -5.

Solved.

---------------

See the lesson
    - Solution of the linear system of two equations in two unknowns by the Elimination method
in this site.